AS = 362.0 J/K AS = -107.0 J/K Given the information A+B 2D AH = -721.2 kJ CD AH* = 577.0 kJ calculate AG at 298 K for the
Answer #1

Dues Solve Given, A +B-72D OH°= -721.2K) ; Os = 3625/K -0. Aho = 577 KJ; os= 107] lk. -6 C AD CS Scanned with CamScanner] Using; ch=blo -TOS where; 66°= Gribbs energy; UH°= Enthalpy: os - Entropy For equation I; DAYE -721.28 103 298KX 362] 1k 82